What is the solution to 5x+7y=-23 and -3x+y=19. (type an ordered pair)substitution method or elimination?? repost if this is not what you inquired...be more specific please.
Answer by haileytucki(390)

5x+7y=-23_-3x+y=19
Since -3x does not contain the variable to solve for, move it to the right-hand side of the equation by adding 3x to both sides.
5x+7y=-23_y=3x+19
Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is 3x+19.
5x+7(3x+19)=-23_y=3x+19
Multiply 7 by each term inside the parentheses.
5x+21x+133=-23_y=3x+19
Since 5x and 21x are like terms, add 21x to 5x to get 26x.
26x+133=-23_y=3x+19
Since 133 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 133 from both sides.
26x=-133-23_y=3x+19
Subtract 23 from -133 to get -156.
26x=-156_y=3x+19
Divide each term in the equation by 26.
(26x)/(26)=-(156)/(26)_y=3x+19
Simplify the left-hand side of the equation by canceling the common factors.
x=-(156)/(26)_y=3x+19
Simplify the right-hand side of the equation by simplifying each term.
x=-6_y=3x+19
Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is -6.
x=-6_y=3(-6)+19
Multiply 3 by each term inside the parentheses.
x=-6_y=-18+19
Add 19 to -18 to get 1.
x=-6_y=1
This is the solution to the system of equations.
x=-6_y=1